A previous survey shows that some 60% of customers prefer brand X. A sample survey is made of 75 customers, what is the probability that more than 70% of the customers do indeed prefer brand X?
Answer by stanbon(75887)

A previous survey shows that some 60% of customers prefer brand X. A sample survey is made of 75 customers, what is the probability that more than 70% of the customers do indeed prefer brand X?
--------
z(0.7) = (0.7-0.6)/sqrt[0.6*0.4/75] = 1.77
-------------
P(p > 0.7) = P(z > 1.77) = 0.0385
=================
Cheers,
Stan H.
=================